OpenStudy (anonymous):

Find the equation of the line through points (-8, 5) and (-8, 8).

OpenStudy (anonymous):

x = -8

OpenStudy (anonymous):

equation of line = y=mx+b here m = slope m= (y2-y1)/(x2-x1) = (8-5)/(-8-8) = -3/16 substitute any coordinate in the equation of line(y=mx+b), and here in substituting (-8, 5) 5= (-3/16 )*-8 +b 5= 3/2 +b ; solving for b b= 5-3/2 = (10-3)/2= 7/2 there for equation of line is y= -3/16 x + 7/2

OpenStudy (anonymous):

@ anetika Mathematica 8 says that the point (-8,8) does not lie on the following line:\[y=\frac{7}{2}-\frac{3 x}{16} \]\[ 8 \neq 5 \]